OpenStudy (calculusxy):

MEDAL WILL BE GIVEN AND FAN!! Jana has a rectangular sheet of cardboard that measures 10 inches by 18 inches. What is the maximum number of 3-inch by 3-inch squares that she can cut from this sheet of cardboard? (A) 15 (B) 17 (C) 18 (D) 19 (E) 20

OpenStudy (anonymous):

Well find area of both and divide.. 10*18=180... 3*3=9... 180/9=20 SO its E

OpenStudy (calculusxy):

@Gerardo_cast23 However, they say that it is C. Can you help?